# On the approximability of random-hypergraph MAX-3-XORSAT problems with quantum algorithms

### v1

@@ V.2 Performance of greedy classical algorithms
To explore the classical difficulty of our PPSPs, we applied a greedy local search algorithm adapted from Bellitti et al. (2021). This algorithm is straightforward; we start with a random bitstring. Then, beginning at each step, we calculate $k=$ ”number of unsatisfied constraints minus the number of satisfied constraints” associated with each bit. We then calculate the fraction of bits, $f_{k}$, belonging to each $k$ value. Note that we only care when $k>0$ because these are the cases where flipping a particular bit will lower the energy. Using some weight function, $w(k)$, we select a $k$ value with normalized probabilities $\propto w(k)*f_{k}$ and flip a bit with that $k$ value. If there are multiple bits belonging to the $k$ value chosen, we choose a bit in this set with uniform probability to flip. This is repeated until the configuration finds itself in a minimum, and the algorithm halts.
@@
Figure 4: Per-shot probability of finding the true ground state with the quasi-greedy classical algorithm defined in section V.2 as adapted from Bellitti et al. (2021), for the four constraint densities studied in this work and planted solution unsatisfied fraction $\epsilon=0.1$. The probability decays superpolynomially with $N$; fits are to $a2^{-bN}$ for $N_{C}=\left\{2,4,6\right\}N$ and $a2^{-b\sqrt{N}}$ for $N_{C}=1.5N^{3/2}$; the scaling fairly well matches the empirically observed Eq. (34). Each datapoint is the average of $10^{5}$ shots. As the constraint density increases, the basin of attraction widens, and the problem becomes easier–though still exponentially scaling–for local update classical routines. Each shot consists of $O\left(N\right)$ local updates so the time to solution scales essentially as the inverse of this probability.
This expected inefficiency of simulated annealing/parallel tempering for this problem can easily be inferred from the results plotted in Figs. 4 and 5. Namely, in all cases the algorithm will find a single relatively deep minimum with high probability at each shot, leading to a super-polynomial cost to escape from it in algorithms simulating a thermal bath. Interestingly, as $N_{C}/N$ increases for fixed unsatisfied ground state fraction $\epsilon$, we find that the decay exponent of the per-shot probability of finding the ground state, $P_{\mathrm{GS}}\left(N\right)$, monotonically decreases, suggesting that the basin of attraction of the true ground state is widening as the problem becomes more extremal. In fact, for $N_{C}\geq 2N$, we empirically observe that the per-shot probability of finding the planted ground state has the approximate scaling
@@
As shown in Fig. 5, these construction rules yield in a set of instances which are hard to approximate in practice. If we let $N_{C}/N$ grow slowly with $N$, e.g. as $\ln\left(N\right)$ or $\sqrt{N}$, then as $N\to\infty$, the probability of finding any states with energies any $O\left(1\right)$ fraction better than random guessing, decays superpolynomially in $N$. And since the unsatisfied fraction $\epsilon$ in the ground state is small but nonzero, Gaussian elimination cannot be used to efficiently find the solution, forcing classical computers to rely on local update algorithms stymied by entropic barriers. It is of course possible that some clever algorithm could be written to exploit our PPSP structure to efficiently solve or approximate these instances classically; we merely claim hardness for generic methods based on local updates. Our PPSP construction rules can easily be generalized to other CSPs, and we suggest that they could prove to be a useful tool for exploring practical approximation hardness in other contexts.
@@
Figure 5: Classical approximation hardness of PPSPs using the quasi-greedy classical algorithm, for constraint densities (clockwise from top left) $N_{C}=\left\{2N,4N,6N,1.5N^{3/2}\right\}$. In each figure we plot the probability that a given shot returns an energy below $qE_{\mathrm{GS}}$ for various choices of $q$. For all the fixed $N_{C}/N$ problem classes we find an empirical approximation threshold $q=A_{g}$ below which finding states becomes superpolynomially hard, and that this value decreases as $N_{C}/N$ increases. For the case $N_{C}=1.5N^{3/2}$ (bottom left), this value steadily drops, as discussed in the text.
V.3 Performance of high-depth QAOA for this problem
@@
Figure 6: Performance of QAOA as an approximator for $N_{C}=\left\{2,4,6,3\sqrt{N}/2\right\}N$ (clockwise from top left), with problem and algorithm parameters described in-text. Plotted are the probabilities $P_{q}\left(N\right)=P\left(E\leq qE_{\mathrm{GS}}\right)$ for $q$ running from 0.5 to 0.85 in steps of 0.05 (top panels) and 0.25 to 0.6 (bottom panels). The results for the larger constraint densities thus plot a weaker approximation range. Thick straight lines correspond to simple exponential fits where $P_{q}\left(N\right)$ is not decaying, dashed straight lines correspond to exponential decay, and thin lines between points are included for visual clarity. As summarized in table 1, these results do not represent a meaningful improvement over the classical greedy result (Fig. 5), though in some cases where $P_{q}\left(N\right)$ decays exponentially for both approaches, the exponent for QAOA may be better. These results were obtained using the Fujitsu Quantum Simulator, a classical HPC system.
@@
To make firm points of comparison, alongside the simulations of spectrally folded optimization itself we extensively benchmarked high-depth QAOA and quasi-greedy Bellitti et al. (2021); Kowalsky et al. (2021) classical algorithms on these instances. For our high-depth QAOA simulations, we formulated our algorithm to mimic trotterized time evolution over a total time $t_{f}$, with:
@@
In all the presented data we used $t_{f}=N/32$ and $dt\simeq 0.05$. Individual shots use a random evolution time between $32t_{f}/$ 3 and $34t_{f}/$ 3; this runtime averaging produces substantially more reliable scaling, particularly when probabilities are small. These parameters were chosen by trial and error for smaller systems; we observed that the probabilities of finding the ground state and other low-energy states increased sublinearly with $t_{f}$ beyond this point. The relative improvements of the probabilities of returning low-lying states were similarly sublinear. No sophisticated numerical or iterative optimization methods were used, for the reasons discussed in section II.5.
@@
The results of our QAOA simulations, of 1000 random PPSPs for each choice of $N$ running from 8 to 30, are shown in Fig. 6. The probability of finding the ground state, shown in Fig. 11, decays exponentially with an exponent very close to that in Eq. (15), which we find remarkable given the simplifying assumptions in that derivation, and that it does not use the more sophisticated techniques used to compute tunneling rates between semiclassical minima. Further, this exponent displays only small variations with constraint density and is nearly identical in all four cases. Smaller system studies for other constraint densities all yielded very similar results for $P_{\mathrm{GS}}$, as predicted by Eq. (15).
@@
Turning to approximation hardness, being relatively sparse, the $N_{C}=2N$ problems are fairly well-approximated by QAOA, with the algorithm returning strings within $q=0.75$ with constant or saturating probability; we attribute this to the presence of many competing minima with energies not far from $E_{\mathrm{GS}}$. In contrast, for $N_{C}=4N$ the algorithm’s performance for approximation degrades, with clear exponential decay for approximation ratios better than $q=0.55$. For higher constraint densities approximation becomes even more difficult, decaying exponentially below $q=0.45$ for $N_{C}=6N$ and 0.25 for $N_{C}=3N^{3/2}/2$. We expect decay at sufficiently large $N$ for any constant fraction in that case, but cannot simulate larger system sizes. Crucially, the thresholds $q_{a}$ we measure are nearly identical to those found by the classical greedy algorithm (figure 5), and no signatures of an exponential quantum advantage in these instances can be seen.
@@
Interestingly, as $N_{C}/N$ increases, the probabilities of finding states comparatively close to $\left|G\right>$ in Hamming distance improve (see Fig. 11), but the probabilities of finding states close in energy worsen. We attribute this behavior to the high energy clustering phase conjectured in Sec. V.2. Empirically for our PPSPs there is a high density of local minima with energies $E\geq q_{a}E_{\mathrm{GS}}$, and if $q_{a}$ is relatively close to 1 it becomes harder for high-depth QAOA to find local excitations near the planted ground state, as the probability amplitudes will be spread over increasingly many competing minima and their own basins of attraction. Conversely, as $q_{a}$ decreases with increasing $N_{C}/N$, the probability of finding local excitations relatively near $\left|G\right>$ increases though still decays exponentially, as the low energy minima far from $\left|G\right>$ are at proportionally higher energies and thus do not compete directly with few-flip states. That the thresholds $q_{a}$ for QAOA match those of the greedy classical algorithm further supports the interpretation of an energy threshold above which local minima become common.
@@
Comparing the classical and established quantum methods, we find that, for these approximation hard instances, QAOA performs very poorly for finding the ground state but less poorly for approximation below the classical hardness threshold $q_{a}$ (see Table 1), with decay exponents that are much closer to the classical result. In some higher constraint density cases our fits produced favorable exponents for approximation with QAOA but our range of $N$ here is smaller than we would prefer to claim any relative quantum advantage absent theoretical justification. Nonetheless, both methods show clear super-polynomial decay per-shot for approximate optimization below the energy range where local minima are dense. With these results in hand, we now turn to folded quantum optimization, which maintains an approximation guarantee regardless of the problem’s constraint density.
VI Numerical tests of spectral folding variations
@@
Having numerically confirmed the expectation that our PPSPs are superpolynomially hard for classical and prior quantum approaches, for both exact and approximate optimization, we now present the results of our folded quantum optimization simulations. We simulated both the trial minimum annealing and interpolation (e.g. AQC) variations. In all cases we chose runtimes increased linearly in $N$, albeit with larger prefactors than in the QAOA simulations (which used $t_{f}=N/32$). A longer runtime further helps reduce the potential influence of diabatic local heating as the Hamiltonian parameters are varied. Run and ramp times that are too short can lead to artificially poor scaling, arising from the formation of local excitations as the transverse field is turned on or off too quickly. This is fundamentally a different, and much more prosaic, issue, than decaying collective tunneling rates, but can be difficult to distinguish when our only measures are energy and Hamming distance from $\left|G\right>$.
@@ VI.1 Spectrally folded adiabatic interpolation performance
We first present simulations of the AQC variation of spectral folding, in figures 7 and 12. For the AQC variation, we followed the procedure in Eq. (35), with a quadratic $H_{\mathrm{fold}}$ (Eq. 5) in place of $H_{P}$ and $f\left(t\right)=\left(1-t/t_{f}\right)^{1/4}$ instead of $\sqrt{1-t/t_{f}}$, with an average runtime $t_{f}=N/24$. This schedule modification was found to improve scaling at higher approximation ratios. In all cases individual shots are runtime averaged between $32t_{f}/$ 3 and $34t_{f}/$ 3 as in our QAOA simulations.
All these parameter choices are the result of intuition, trial and error, and a desire for simulations to complete in reasonable amounts of time; none are particularly optimal. Nonetheless, as shown in figure 7, in all but one case we were able to meet the approximation target $A=0.75$, but not exceed it; for $N_{C}=4N$. Our spectral folding methods are also much better at reliably returning states close to $G$ in Hamming distance, consistent with the approximation guarantee (see Fig. 12). We likewise tested increasing $A$ to 0.85 in simulations up to $N=24$ with this variation, and found improved prefactors but no improvement in scaling (see Fig. 14). This suggests that we have found the performance ceiling for this approach. Interestingly, the worst case $q_{a}=0.7$ observed for this method is very close to the approximation ratio of $\sim 0.68$ predicted in section IV.5, using a more simplified analysis than was employed for the TMA variation.

### v3

@@ V.2 Performance of quasi-greedy classical algorithms
To explore the classical difficulty of our PPSPs, we applied a greedy local search algorithm adapted from Bellitti et al. (2021). This algorithm is straightforward; we start with a random bitstring. Then, beginning at each step, we calculate $k=N_{\rm{sat},j}-N_{\rm{unsat},j}$ associated with each bit $j$. We then calculate the fraction of bits, $f_{k}$, belonging to each $k$ value. Note that we only care when $k>0$ because these are the cases where flipping a particular bit will lower the energy. Using some weight function, $w(k)$, we select a $k$ value with normalized probabilities $\propto w(k)*f_{k}$ and flip a bit with that $k$ value. If there are multiple bits belonging to the $k$ value chosen, we choose a bit in this set with uniform probability to flip. This is repeated until the configuration finds itself in a minimum, and the algorithm halts.
@@
Figure 6: Per-shot probability of finding the true ground state with the quasi-greedy classical algorithm defined in section V.2 as adapted from Bellitti et al. (2021), for the four constraint densities studied in this work and planted solution unsatisfied fraction $\epsilon=0.1$. The probability decays superpolynomially with $N$; fits are to $a2^{-bN}$ for $N_{C}=\left\{2,4,6\right\}N$ and $a2^{-b\sqrt{N}}$ for $N_{C}=1.5N^{3/2}$; the scaling matches the empirically observed Eq. (43) fairly well. Each datapoint is the average of $10^{5}$ shots. As the constraint density increases, the basin of attraction widens, and the problem becomes easier–though still exponentially scaling–for local update classical routines. Each shot consists of $O\left(N\right)$ local updates so the time to solution scales essentially as the inverse of this probability.
This expected inefficiency of simulated annealing/parallel tempering for this problem can easily be inferred from the results plotted in Figs. 6 and 7. Namely, in all cases the algorithm will find a single relatively deep minimum with high probability at each shot, leading to a super-polynomial cost to escape from it in algorithms simulating a thermal bath. Interestingly, as $N_{C}/N$ increases for fixed unsatisfied ground state fraction $\epsilon$, we find that the decay exponent of the per-shot probability of finding the ground state, $P_{\mathrm{GS}}\left(N\right)$, monotonically decreases, suggesting that the basin of attraction of the true ground state is widening as the problem becomes more extremal. In fact, for $N_{C}\geq 2N$, we empirically observe that the per-shot probability of finding the planted ground state has the approximate scaling
@@
As shown in Fig. 7, these construction rules yield in a set of instances which are hard to approximate in practice. If we let $N_{C}/N$ grow slowly with $N$, e.g. as $\ln\left(N\right)$ or $\sqrt{N}$, then as $N\to\infty$, the probability of finding any states with energies any $O\left(1\right)$ fraction better than random guessing, decays superpolynomially [^12] We note again the recent result of d’Orsi and Trevisan (2022), who showed that for $N_{C}/N>O\left(N^{1/2}\right)$ the problem can be efficiently approximated classically; our PPSPs with $N_{C}/N=1.5N^{1/2}$ are close to this threshold but do not cross it. in $N$. And since the unsatisfied fraction $\epsilon$ in the ground state is small but nonzero, Gaussian elimination cannot be used to efficiently find the solution, forcing classical computers to rely on local update algorithms stymied by entropic barriers. It is of course possible that some clever algorithm could be written to exploit our PPSP structure to efficiently solve or approximate these instances classically; we merely claim hardness for generic methods based on local updates. Our PPSP construction rules can easily be generalized to other CSPs, and we suggest that they could prove to be a useful tool for exploring practical approximation hardness in other contexts.
@@
Figure 7: Classical approximation hardness of PPSPs using the quasi-greedy classical algorithm, for constraint densities (clockwise from top left) $N_{C}=\left\{2N,4N,6N,1.5N^{3/2}\right\}$. In each figure we plot the probability that a given shot returns an energy below $qE_{\mathrm{GS}}$ for various choices of $q$ (note varying scales on probability $y$-axis). For all the fixed $N_{C}/N$ problem classes we find an empirical approximation threshold $q=A_{g}$ below which finding states becomes superpolynomially hard, and that this value decreases as $N_{C}/N$ increases. For the case $N_{C}=1.5N^{3/2}$ (bottom left), this value steadily drops, as discussed in the text.
V.3 Performance of high-depth TAQC for this problem
@@
Figure 8: Performance of TAQC as an approximator for $N_{C}=\left\{2,4,6,3\sqrt{N}/2\right\}N$ (clockwise from top left), with problem and algorithm parameters described in section V.3. Plotted are the probabilities $P_{q}\left(N\right)=P\left(E\leq qE_{\mathrm{GS}}\right)$ for $q$ running from 0.5 to 0.85 in steps of 0.05 (top panels) and 0.25 to 0.6 (bottom panels). The results for the larger constraint densities thus plot a weaker approximation range. Thick straight lines correspond to simple exponential fits where $P_{q}\left(N\right)$ is not decaying; dashed straight lines correspond to exponential decay; jagged thin curves are a guide to the eye; only points represent actual data. As summarized in Table 1, these results do not represent a meaningful improvement over the classical greedy result (Fig. 7), though in some cases where $P_{q}\left(N\right)$ decays exponentially for both approaches, the exponent for TAQC may be better. These results were obtained using the Fujitsu Quantum Simulator, a classical HPC system.
Figure 9: Probabilities of returning states within Hamming distance $D_{H}=\left\{2/5,1/3,1/4,1/5,1/8,0\right\}N$ flips from the ground state $G$ for QAOA (top to bottom curves), for $N_{C}=\left\{2,4,6,3\sqrt{N}/2\right\}N$ (clockwise from top left) and application parameters in section V.3. As seen in the figures, the per-shot probability of finding the ground state is essentially independent of constraint density and tracks the prediction $N2^{-0.28N}$ in Eq. 25. On the other hand, the probabilities of returning states at various extensive fractional Hamming distances, e.g. $N/4$ or fewer flips, decay much more slowly, owing to the exponentially large number of target states, and the task of finding states comparatively close to $G$ becomes easier as $N_{C}/N$ increases. That said, these probabilities all decay exponentially with $N$ in all cases, consistent with the worsening approximation ratios observed in Figure 8. In particular, as discussed in section V.3, these results allow us to test the prediction in section II.3 that the per-state matrix elements to mix with excited states can collapse as the energy increases. The numbers adjacent to each curve on the plots are the best fit for the decay exponent $b$ in $\Omega_{0j}^{2}\right>_{j}\propto 2^{-bN}$; as the Hamming distance (and thus average energy, per Eq. 18) increases, these matrix elements decay much more rapidly, exceeding the exponential growth of the number of target states and confirming our expectation that AQC is a poor approximation algorithm in the worst case.
@@
To make firm points of comparison, alongside the simulations of spectrally folded optimization itself we extensively benchmarked high-depth TAQC on these instances. For our high-depth TAQC simulations, we formulated our algorithm to approximate continuous time evolution over a total time $t_{F}$, with:
@@
In all the presented data we used $t_{F}=N/32$ and $dt\simeq 0.05$. Individual shots use a random evolution time between $32t_{F}/$ 3 and $34t_{F}/$ 3; this runtime averaging produces substantially more reliable scaling, particularly when probabilities are small. These parameters were chosen by trial and error for smaller systems; we observed that the probabilities of finding the ground state and other low-energy states increased sublinearly with $t_{F}$ beyond this point. The relative improvements of the probabilities of returning low-lying states were similarly sublinear. As discussed in section II.5, we focused TAQC rather than QAOA to avoid the overhead of angle optimization and to have easily interpretable results. No numerical or iterative optimization methods were used, nor was any per-instance tuning employed, for this algorithm or any other algorithm in this work.
@@
The results of our TAQC simulations, of 1000 random PPSPs for each choice of $N$ running from 8 to 30, are shown in Fig. 8. The probability of finding the ground state, shown in Fig. 9, decays exponentially with an exponent very close to that in Eq. (25), which we find remarkable given the simplifying assumptions in that derivation, and that it does not use the more sophisticated techniques used to compute tunneling rates between semiclassical minima. Further, this exponent displays only small variations with constraint density and is nearly identical in all four cases. Smaller system studies for other constraint densities all yielded very similar results for $P_{\mathrm{GS}}$, as predicted by Eq. (25).
@@
Turning to approximation hardness, being relatively sparse, the $N_{C}=2N$ problems are fairly well-approximated by TAQC, with the algorithm returning strings within $q=0.75$ with constant or saturating probability; we attribute this to the presence of many competing minima with energies not far from $E_{\mathrm{GS}}$. In contrast, for $N_{C}=4N$ the algorithm’s performance for approximation degrades, with clear exponential decay for approximation ratios better than $q=0.55$. For higher constraint densities approximation becomes even more difficult, decaying exponentially below $q=0.45$ for $N_{C}=6N$ and 0.25 for $N_{C}=3N^{3/2}/2$. We expect decay at sufficiently large $N$ for any constant fraction in that case, but cannot simulate larger system sizes. Crucially, the thresholds $q_{a}$ we measure are nearly identical to those found by the classical greedy algorithm (figure 7), and no signatures of an exponential quantum advantage in these instances can be seen.
@@
Interestingly, as $N_{C}/N$ increases, the probabilities of finding states comparatively close to $\left|G\right>$ in Hamming distance improve (see Fig. 9), but the probabilities of finding states close in energy worsen. We attribute this behavior to the high energy clustering phase conjectured in Secs. II.4 and V.2. Empirically for our PPSPs there is a high density of local minima with energies $E\geq q_{a}E_{\mathrm{GS}}$, and if $q_{a}$ is relatively close to 1 it becomes harder for high-depth TAQC to find local excitations near the planted ground state, as the probability amplitudes will be spread over increasingly many competing minima and their own basins of attraction. Conversely, as $q_{a}$ decreases with increasing $N_{C}/N$, the probability of finding local excitations relatively near $\left|G\right>$ increases though still decays exponentially, as the low energy minima far from $\left|G\right>$ are at proportionally higher energies and thus do not compete directly with few-flip states. That the thresholds $q_{a}$ for TAQC match those of the greedy classical algorithm further supports the interpretation of an energy threshold above which local minima become common.
Further, these results allow us to directly test the qualitative predition in section II.3 that the matrix elements for the initial paramagnetic state to mix with excited states can collapse as the energy increases, due to the weakening transverse field and destructive interference. As shown in FIG. 9, the probabilities of finding states within fractional Hamming distances $\left\{2/5,1/3,1/4,1/5,1/8\right\}N$ from $G$ all decay exponentially, albeit with much smaller exponents than the probability of finding $G$. We emphasize that these smaller exponents mask the fact that the per-state mixing rate $\Omega_{0j}^{2}\right>_{j}$ is decaying extremely rapidly, since the total success probability for finding states within $D_{H}$ flips is given by
$P_{D_{H}}\left(t_{F}\right)\propto\frac{\left<\Omega_{0j}^{2}\right>_{j}t_{F}}{W}N_{D_{H}},$ (45)
where $N_{D_{H}}$ is the number of states at this distance and is given by Eq. 40. Since $N_{D_{H}}$ grows exponentially quickly, for the total $P_{D_{H}}$ to decay exponentially $\Omega_{0j}^{2}\right>_{j}$ must decay rapidly, and the resulting fit exponents in FIG. 9 are stark evidence of this. In section II.3, we identified weakening fields and destructive interference as good reasons to not be confident in TAQC as a potent approximation algorithm, even though we were not able to perform a rigorous analytical calculation due to the complexity involved. Our numerical results confirm this suspicion. Spectrally filtered optimization was originally formulated to resolve these two issues, and as shown in the next section, leads to dramatically different behavior and approximation performance.
@@
Comparing the classical and established quantum methods, we find that, for these approximation hard instances, TAQC performs very poorly for finding the ground state but less poorly for approximation below the classical hardness threshold $q_{a}$ (see Table 1), with decay exponents that are much closer to the classical result. In some higher constraint density cases our fits produced favorable exponents for approximation with TAQC but our range of $N$ here is smaller than we would prefer to claim any relative quantum advantage absent rigorous theoretical justification. Nonetheless, both methods show clear super-polynomial decay per-shot for approximate optimization below the energy range where local minima are dense. With these results in hand, we now turn to folded quantum optimization, which maintains an approximation guarantee regardless of the problem’s constraint density.
VI Numerical tests of spectrally filtered quantum optimization
@@
Having numerically confirmed the expectation that our PPSPs are superpolynomially hard for classical and prior quantum approaches, for both exact and approximate optimization, we now present the results of our folded quantum optimization simulations. We simulated both the trial minimum annealing and interpolation (e.g. AQC) variations. In all cases aside from MSFO, we chose runtimes increased linearly in $N$, albeit with larger prefactors than in the TAQC simulations (which used $t_{F}=N/32$). Following the discussion in section II.5, a longer runtime further helps reduce the potential influence of diabatic local heating as the Hamiltonian parameters are varied. Run and ramp times that are too short can lead to artificially poor scaling, arising from the formation of local excitations as the transverse field is turned on or off too quickly. This is fundamentally a different, and much more prosaic, issue, than decaying collective tunneling rates, but can be difficult to distinguish when our only measures are energy and Hamming distance from $\left|G\right>$.
@@ VI.1 Spectrally folded adiabatic interpolation performance
We first present simulations of the AQC variation of spectral folding, in figures 10 and 16. For the AQC variation, we followed the procedure in Eq. (44), with a quadratic $H_{\mathrm{fold}}$ (Eq. 8, without the symmetrization at high energies) in place of $H_{P}$ and $f\left(t\right)=\left(1-t/t_{F}\right)^{1/4}$ instead of $\sqrt{1-t/t_{F}}$, with an average runtime $t_{F}=N/24$ and $dt=0.0325$. This schedule modification was found to improve scaling at higher approximation ratios. In all cases individual shots are runtime averaged between $32t_{F}/$ 3 and $34t_{F}/$ 3 as in our TAQC simulations.
As shown in figure 10, in all but one case we were able to meet the approximation target $A=0.75$, but not exceed it; for $N_{C}=4N$ we were limited to $q_{a}\simeq 0.7$. Our spectral folding methods are also much better at reliably returning states close to $G$ in Hamming distance, consistent with the approximation guarantee (see Fig. 16). We likewise tested increasing $A$ to 0.85 in simulations up to $N=24$ with this variation, and found improved prefactors but no improvement in scaling (see Fig. 18). This suggests that we have found the performance ceiling for this approach. Interestingly, the worst case $q_{a}=0.7$ observed for this method is not far below the approximation ratio of $\sim 0.8$ predicted in section IV.3.3, using a much more simplified analysis than was employed for the TMA and MSFO variations.